# White dwarf eccentricity fluctuation and dissipation by AGB convection

### v1

@@ Abstract
Millisecond pulsars with white dwarf companions have typical eccentricities $e\sim 10^{-6}-10^{-3}$. The eccentricities of helium white dwarfs are explained well by applying the fluctuation–dissipation theorem to convective eddies in their red giant progenitors. We extend this theory to more massive carbon–oxygen (CO) white dwarfs with asymptotic giant branch (AGB) progenitors. Due to the radiation pressure in AGB stars, the dominant factor in determining the remnant white dwarf’s eccentricity is the critical residual hydrogen envelope mass $m_{\rm env}$ required to inflate the star to giant proportions. Using a suite of mesa stellar evolution simulations with $\Delta m_{\rm c}=10^{-3}\,{\rm M}_{\sun}$ core-mass intervals, we resolved the AGB thermal pulses and found that the critical $m_{\rm env}\propto m_{\rm c}^{-6}$. This steep dependence causes the $e(m_{\rm c})$ relation to turn over, such that $e\sim 3\times 10^{-3}$ almost independently of the remnant CO white dwarf’s mass $m_{\rm c}$. Nearly all of the measured eccentricities lie below this robust theoretical limit, indicating that the eccentricity is damped during the common-envelope inspiral that follows the unstable Roche-lobe overflow of the AGB star. Specifically, we focused on white dwarfs with median masses $m_{\rm c}>0.6\,{\rm M}_{\sun}$. These massive white dwarfs begin their inspiral with practically identical orbital periods and eccentricities, eliminating any dependence on the initial conditions. For this sub-sample, we find an empirical relation $e\propto P^{3/2}$ between the final period and eccentricity that is much tighter than previous studies – motivating theoretical work on the eccentricity evolution during the common envelope phase.
@@
pubyear: 2024pagerange: White dwarf eccentricity fluctuation and dissipation by AGB convectionWhite dwarf eccentricity fluctuation and dissipation by AGB convection
@@ 1 Introduction
The remainder of this paper is organized as follows. In Section 2 we analyse the structure of RGB and AGB giant stars during mass transfer. In Section 3 we calculate the evolution of the orbital period and eccentricity during and after Roche-lobe overflow. We compare the theory to the observed periods and eccentricities of binary millisecond pulsars in Section 4. We summarize and discuss our results in Section 5.
@@ 2.2 Numerical calculations
Our numerically computed $m_{\rm env}^{\rm crit}$ decreases steeply with increasing $m_{\rm c}$ (Fig. 3). [^2] Extrapolating our power-law fit reveals that $m_{\rm env}^{\rm crit}=m_{\rm c}$ for $m_{0}\approx 0.09\,{\rm M}_{\sun}$, which is close to the hydrogen burning limit separating main-sequence stars from degenerate objects. Plugging $m_{0}$ in equation (19) also roughly reproduces the main-sequence/degenerate radius at this mass $r_{0}\approx 0.1\,{\rm R}_{\sun}$. This is not a coincidence: for $m_{0}$, the temperature above the degenerate core $T_{\rm c}$ is equal to the hydrogen burning temperature in main-sequence stars, such that the radius $r_{\rm c}$, density $\rho_{\rm c}$, and luminosity $L$ of the burning shell are similar to those of a main-sequence star at this mass. In other words, $r_{\rm env}\sim r_{\rm c}\sim r_{0}$ and $m_{\rm sh}\sim m_{\rm c}\sim m_{0}$ – providing a useful normalization point for red giants. While this qualitative behaviour is predicted by the analytical equations (8) and (18), our fitted power-laws differ considerably from equation (8). Possible sources for this discrepancy include the variation of the nuclear burning exponent $\nu$ with temperature (and therefore with $m_{\rm c}$), the logarithmic factor in equation (18), which also depends on $m_{\rm c}$, and the marginal degeneracy of the hot core, which leads to deviations from $r_{\rm c}\propto m_{\rm c}^{-1/3}$ (Refsdal & Weigert, 1970). In addition, our single shell analytical theory (Section 2.1) does not capture the alternating burning of the double shell structure on the AGB and the resulting thermal pulses (e.g. Belloni et al., 2024b).
@@ 3 Orbital evolution
The tidal circularization time-scale on which this eccentricity equilibrium is established scales as $(r_{\rm env}/R)^{-8}$ (Zahn, 1977; Phinney, 1992; Ginzburg & Chiang, 2022). [^3] The same convective eddies that perturb the orbit and excite the eccentricity also damp the eccentricity by dissipating the energy stored in the tidal bulges that the pulsar raises. This is a manifestation of the fluctuation–dissipation theorem (Phinney, 1992). As long as the giant star fills its Roche lobe such that $r_{\rm env}/R\sim(m_{\rm c}/M)^{1/3}$, this time-scale is short enough to enforce equation (25) as the star evolves. However, once the envelope’s mass is reduced to $m_{\rm env}\approx m_{\rm env}^{\rm crit}$ and $r_{\rm env}$ contracts by a factor of a few, the circularization time-scale exceeds the contraction time-scale. At this point, the contraction ‘runs away’ and tides are no longer able to adjust the eccentricity. As a result, the eccentricity freezes when $m_{\rm env}\approx m_{\rm env}^{\rm crit}$ and $r_{\rm env}\sim r_{\rm env}^{\rm max}$. By plugging $r_{\rm env}$ from equation (17) in equation (25) we find that the eccentricity scales approximately as
@@ 4 Observations
Figure 4: Observed field binary millisecond pulsars from the ATNF Pulsar Catalogue with helium (squares) and CO/ONe (blue circles) white dwarf companions. Top panel: unlike helium white dwarfs, the orbital periods of CO white dwarfs are inconsistent with the Roche-lobe overflow of their progenitors (solid line), with a radius $r_{\rm env}$ given by equation (19). The dashed line marks the detailed evolutionary calculations of Tauris & Savonije (1999). Bottom panel: the eccentricities of most helium white dwarfs are set by convection in their progenitors’ envelopes when they detach from their Roche lobe (Phinney 1992; the exceptional eMSPs are coloured in red). We extend this theory to CO white dwarfs using equation (25), with envelope masses $m_{\rm env}$ fitted from Fig. 3. This limit (solid line) provides the initial conditions for the subsequent inspiral from Roche-lobe overflow to the current periods.
@@
The observed orbital periods of CO white dwarfs, on the other hand, are orders of magnitude shorter than the periods at which their progenitors filled their Roche lobe – indicating an inspiral following an unstable mass transfer episode (i.e. a common envelope event). We estimate the eccentricity at the onset of the inspiral by extending the Phinney (1992) model using equation (25): by substituting $r_{\rm env}$ from equation (19) and $m_{\rm env}$ from the analytical fits in Fig. 3, we derive a more accurate version of equation (27), which is given by the solid line in the bottom panel of Fig. 4. We normalize our $e(m_{\rm c})$ relation, as well as the $e\propto m_{\rm c}^{6}$ dashed line, to coincide with Phinney (1992) at an orbital period of 100 d ($m_{\rm c}\approx 0.3\,{\rm M}_{\sun}$). We emphasize that this normalization is not a freely adjustable fitting parameter, but it was rather calculated theoretically by Phinney (1992) using a more accurate version of our Section 3. As a consequence, our theory has no free parameters, and it implies that convection excites eccentricities $e\sim 3\times 10^{-3}$ for all CO white dwarfs during Roche-lobe overflow, regardless of their mass (because of the turnover in the $e$$m_{\rm c}$ relation).
@@ 5 Summary and discussion
In this paper, we turned our attention to a similar class of binary millisecond pulsars with heavier CO white dwarf companions. The formation history of these systems is less obvious than their helium counterparts (Tauris et al., 2011, 2012). Specifically, their current orbital periods are orders of magnitude shorter than the Roche-lobe filling periods of their presumed AGB progenitor stars, possibly indicating an inspiral within a common envelope following unstable Roche-lobe overflow. Surprisingly, despite this different evolution, the typical measured eccentricities of these two classes are rather similar ($e\sim 10^{-6}-10^{-3}$), leading us to speculate whether they were excited by the same mechanism. Other models might face a fine-tuning challenge in trying to explain such tiny – but finite – eccentricities.
@@
To answer this question, we extended the Phinney (1992) model to AGB stars that initially fill their Roche lobe. Due to the increasing role of radiation pressure in their interiors, the $P(m_{\rm c})$ relation of these stars flattens (Rappaport et al., 1995), such that the mass of the envelope $m_{\rm env}^{\rm crit}$ – rather than that of the core $m_{\rm c}$ – becomes the dominant factor in determining the eccentricity. By running a suite of mesa simulations with $\Delta m_{\rm c}=10^{-3}\,{\rm M}_{\sun}$ intervals, we resolved the thermal pulses on the AGB and found that $m_{\rm env}^{\rm crit}\propto m_{\rm c}^{-6}$. This steep dependence causes the $e(m_{\rm c})$ relation to turn over, such that AGB convection yields an approximately constant $e\sim 3\times 10^{-3}$, which is almost independent of $m_{\rm c}$.
@@
We compared our theoretical prediction – which has no free parameters – to the observations, finding that almost all CO white dwarf–millisecond pulsar binaries have lower eccentricities. This indicates that orbital eccentricities are damped, rather than excited, during the subsequent common envelope inspiral phase. Quantitatively, we found an empirical relation $e\propto P^{3/2}$ between the final eccentricities and periods of white dwarfs with median masses $m_{\rm c}>0.6\,{\rm M}_{\sun}$. Presumably thanks to the almost uniform Roche-lobe overflow $e$ and $P$ initial conditions in this mass range, our final $e$$P$ relation is much tighter than Hui et al. (2018), who considered less massive CO white dwarfs as well. This observed correlation motivates future theoretical work on the eccentricity evolution during the common envelope phase (see also Dermine et al., 2013; Glanz & Perets, 2021; Szölgyén et al., 2022; Trani et al., 2022).
@@
Although we focused here on radio observations of white dwarfs around millisecond pulsars, our results may be applied to overflowing AGB stars (and their white dwarf remnants) orbiting other types of stars as well. Specifically, the Gaia mission has recently uncovered a population of white dwarf–main sequence wide binaries that could have formed similarly (Shahaf et al., 2024; Yamaguchi et al., 2024b). Another example is the class of self-lensing binaries found by the Kepler spacecraft, which lie relatively close to (and below) the Roche-lobe filling $P(m_{\rm c})$ line of AGB stars (Kruse & Agol, 2014; Kawahara et al., 2018; Belloni et al., 2024a; Yamaguchi et al., 2024a).

### v2

@@ Abstract
Millisecond pulsars with white dwarf companions have typical eccentricities $e\sim 10^{-6}-10^{-3}$. The eccentricities of helium white dwarfs are explained well by applying the fluctuation–dissipation theorem to convective eddies in their red giant progenitors. We extend this theory to more massive carbon–oxygen (CO) white dwarfs with asymptotic giant branch (AGB) progenitors. Due to the radiation pressure in AGB stars, the dominant factor in determining the remnant white dwarf’s eccentricity is the critical residual hydrogen envelope mass $m_{\rm env}$ required to inflate the star to giant proportions. Using a suite of mesa stellar evolution simulations with $\Delta m_{\rm c}=10^{-3}\,{\rm M}_{\sun}$ core-mass intervals, we resolved the AGB thermal pulses and found that the critical $m_{\rm env}\propto m_{\rm c}^{-6}$. The resulting eccentricity $e\sim 3\times 10^{-3}$ is almost independent of the remnant CO white dwarf’s mass $m_{\rm c}$. Nearly all of the measured eccentricities lie below this robust theoretical limit, indicating that the eccentricity is damped during the common-envelope inspiral that follows the unstable Roche-lobe overflow of the AGB star. Specifically, we focused on white dwarfs with median masses $m_{\rm c}>0.6\,{\rm M}_{\sun}$. These massive white dwarfs begin their inspiral with practically identical orbital periods and eccentricities, eliminating any dependence on the initial conditions. For this sub-sample, we find an empirical relation $e\propto P^{3/2}$ between the final period and eccentricity that is much tighter than previous studies – motivating theoretical work on the eccentricity evolution during the common envelope phase. The eccentricities of lower mass CO white dwarfs may be explained by alternative formation channels.
@@
pubyear: 2024pagerange: White dwarf eccentricity fluctuation and dissipation by AGB convection9
@@ 1 Introduction
The remainder of this paper is organized as follows. In Section 2 we analyse the structure of RGB and AGB giant stars during mass transfer. In Section 3 we calculate the evolution of the orbital period and eccentricity during and after Roche-lobe overflow. We compare the theory to the observed periods and eccentricities of binary millisecond pulsars in Section 4. In Section 5 we briefly consider alternative formation paths. We summarize and discuss our results in Section 6.
@@ 2.2 Numerical calculations
Our numerically computed $m_{\rm env}^{\rm crit}$ decreases steeply with increasing $m_{\rm c}$ (Fig. 3). [^2] Extrapolating our power-law fit reveals that $m_{\rm env}^{\rm crit}=m_{\rm c}$ for $m_{0}\approx 0.09\,{\rm M}_{\sun}$, which is close to the hydrogen burning limit separating main-sequence stars from degenerate objects. Plugging $m_{0}$ in equation (19) also roughly reproduces the main-sequence/degenerate radius at this mass $r_{0}\approx 0.1\,{\rm R}_{\sun}$. This is not a coincidence: for $m_{0}$, the temperature above the degenerate core $T_{\rm c}$ is equal to the hydrogen burning temperature in main-sequence stars, such that the radius $r_{\rm c}$, density $\rho_{\rm c}$, and luminosity $L$ of the burning shell are similar to those of a main-sequence star at this mass. In other words, $r_{\rm env}\sim r_{\rm c}\sim r_{0}$ and $m_{\rm sh}\sim m_{\rm c}\sim m_{0}$ – providing a useful normalization point for red giants. While this qualitative behaviour is predicted by the analytical equations (8) and (18), our fitted power-laws differ considerably from equation (8). Possible sources for this discrepancy include the variation of the nuclear burning exponent $\nu$ with temperature (and therefore with $m_{\rm c}$), the logarithmic factor in equation (18), which also depends on $m_{\rm c}$, and the marginal degeneracy of the hot core, which leads to deviations from $r_{\rm c}\propto m_{\rm c}^{-1/3}$ (Refsdal & Weigert, 1970). In addition, our single shell analytical theory (Section 2.1) does not capture the alternating burning of the double shell structure on the AGB and the resulting thermal pulses (e.g. Belloni et al., 2024a).
@@ 3 Orbital evolution
The tidal circularization time-scale on which this eccentricity equilibrium is established scales as $(r_{\rm env}/R)^{-8}$ (Zahn, 1977; Phinney, 1992; Ginzburg & Chiang, 2022). [^3] The same convective eddies that perturb the orbit and excite the eccentricity also damp the eccentricity by dissipating the energy stored in the tidal bulges that the pulsar raises. This is a manifestation of the fluctuation–dissipation theorem (Phinney, 1992). As long as the giant star fills its Roche lobe such that $r_{\rm env}/R\sim(m_{\rm c}/M)^{1/3}$, this time-scale is short enough to enforce equation (25) as the star evolves (we check this in Appendix A). However, once the envelope’s mass is reduced to $m_{\rm env}\approx m_{\rm env}^{\rm crit}$ and $r_{\rm env}$ contracts by a factor of a few, the circularization time-scale exceeds the contraction time-scale. At this point, the contraction ‘runs away’ and tides are no longer able to adjust the eccentricity. As a result, the eccentricity freezes when $m_{\rm env}\approx m_{\rm env}^{\rm crit}$ and $r_{\rm env}\sim r_{\rm env}^{\rm max}$. By plugging $r_{\rm env}$ from equation (17) in equation (25) we find that the eccentricity scales approximately as
@@ 4 Observations
Figure 4: Observed field binary millisecond pulsars from the ATNF Pulsar Catalogue with helium (squares) and CO/ONe (blue circles) white dwarf companions. Top panel: unlike helium white dwarfs, the orbital periods of CO white dwarfs are inconsistent with the Roche-lobe overflow of their progenitors (solid line), with a radius $r_{\rm env}$ given by equation (19). The dashed line marks the detailed evolutionary calculations of Tauris & Savonije (1999). Bottom panel: the eccentricities of most helium white dwarfs are set by convection in their progenitors’ envelopes when they detach from their Roche lobe (Phinney 1992; the exceptional eMSPs are coloured in red). We extend this theory to CO white dwarfs using equation (25), with envelope masses $m_{\rm env}$ fitted from Fig. 3. This limit (solid line) provides the initial conditions for the subsequent inspiral from Roche-lobe overflow to the current periods. The magenta star markers indicate the results of an alternative formation path – Case A/B Roche-lobe overflow – which we discuss in Section 5.
@@
The observed orbital periods of CO white dwarfs, on the other hand, are orders of magnitude shorter than the periods at which their progenitors filled their Roche lobe – indicating an inspiral following an unstable mass transfer episode (i.e. a common envelope event). We estimate the eccentricity at the onset of the inspiral by extending the Phinney (1992) model using equation (25): by substituting $r_{\rm env}$ from equation (19) and $m_{\rm env}$ from the analytical fits in Fig. 3, we derive a more accurate version of equation (27), which is given by the solid line in the bottom panel of Fig. 4. We normalize our $e(m_{\rm c})$ relation, as well as the $e\propto m_{\rm c}^{6}$ dashed line, to coincide with Phinney (1992) at an orbital period of 100 d ($m_{\rm c}\approx 0.3\,{\rm M}_{\sun}$). We emphasize that this normalization is not a freely adjustable fitting parameter, but it was rather calculated theoretically by Phinney (1992) using a more accurate version of our Section 3. As a consequence, our theory has no free parameters, and it implies that convection excites eccentricities $e\sim 3\times 10^{-3}$ for all CO white dwarfs during Roche-lobe overflow, regardless of their mass (because of the nearly flat $e$$m_{\rm c}$ relation at these masses).
@@
5 Alternative formation paths
So far we have focused on ‘Case C’ mass transfer from AGB stars harbouring degenerate CO cores, leading to common envelope inspiral. However, there are alternative paths to form CO white dwarf–millisecond pulsar binaries that avoid both the AGB and the common envelope phases (see Tauris, 2011, for a concise review). For a certain region of the parameter space (i.e. initial donor mass, orbital period, and neutron star mass), mass transfer that is initiated on the main sequence (Case A) or the RGB (Case B) remains stable and the donor ignites its helium core after detaching from the Roche lobe, leaving behind a CO white dwarf remnant (Tauris et al., 2000; Shao & Li, 2012; Misra et al., 2020).
Figure 6: Single stellar evolution of a $4\,{\rm M}_{\sun}$ star. When in a binary, such a star may overflow its Roche lobe while on the main sequence (Case A), RGB (Case B), or AGB (Case C) – depending on the initial orbital period $P$. As a specific example, the magenta star marker indicates the overflow moment for $P=5\,{\rm d}$ around an $M=1.8\,{\rm M}_{\sun}$ neutron star. The binary evolution of the overflowing star in this example is plotted in Fig. 7.
Figure 7: Binary stellar evolution of an initially $4\,{\rm M}_{\sun}$ star on a $P=5\,{\rm d}$ orbit around an $M=1.8\,{\rm M}_{\sun}$ neutron star. After Case B Roche-lobe overflow (see Fig. 6), the star detaches (magenta star markers) and contracts, eventually forming an $m_{\rm c}\approx 0.5\,{\rm M}_{\sun}$ CO white dwarf. This detachment sets the white dwarf’s final orbital period $P$ (top panel) and eccentricity $e$ according to equation (25), with the critical $m_{\rm env}$ inferred from the bottom panel. These results are compared to the observations in Fig. 4.
In Figs 6 and 7 we examine a typical example of such a Case A/B formation scenario: an initially $4\,{\rm M}_{\sun}$ donor with an initial orbital period $P=5\,{\rm d}$ around an $M=1.8\,{\rm M}_{\sun}$ neutron star. We evolved this system using the star_plus_point_mass binary test suite in mesa. Following Case B Roche-lobe overflow, the donor detaches from its Roche lobe at an orbital period $P\approx 35\,{\rm d}$, when its envelope is reduced to a critical mass $m_{\rm env}\approx 6\times 10^{-2}\,{\rm M}_{\sun}$. Similarly to Phinney (1992) and our Section 3, the subsequent contraction of the donor star well below its Roche lobe shuts off tidal interactions and freezes the eccentricity evolution. We apply equation (25) with our regular normalization (which, again, is not a free parameter) and find $e\approx 7\times 10^{-5}$ for the $m_{\rm c}\approx 0.5\,{\rm M}_{\sun}$ CO white dwarf remnant. As expected and verified in Fig. 5, this value is very close to the original Phinney (1992) $e\propto P$ relation, which was derived under similar assumptions; the small difference stems from the relatively weak dependence on $m_{\rm env}/m_{\rm c}$, which was neglected by Phinney (1992).
As seen in Fig. 4, this alternative formation path produces CO white dwarfs with periods and eccentricities that are close to some of the observations. Unlike Case C overflow, our computed $P$ and $e$ (magenta star markers) actually represent the final values of these observables, because the system avoids a common envelope inspiral. Case A/B overflow therefore provides a potential alternative explanation for the measured CO white dwarf eccentricities. This scenario is mainly relevant for lower mass CO white dwarfs (Tauris et al., 2000; Shao & Li, 2012; Misra et al., 2020), i.e. the empty circles in Fig. 5. The higher mass white dwarfs on which we focus in this paper (the filled circles in Fig. 5) probably formed through Case C mass transfer followed by common envelope inspiral, as assumed in the previous sections (see fig. 4 of Tauris, 2011). Our promising results in this section for a typical example of Case A/B evolution motivate a more extensive investigation of the mass, period, and eccentricity ranges of white dwarfs that are formed in this scenario. We defer such a survey of the parameter space to future work, and focus here on the AGB path.
@@ 6 Summary and discussion
In this paper, we turned our attention to a similar class of binary millisecond pulsars with heavier CO white dwarf companions. The formation history of these systems is less obvious than their helium counterparts (Tauris, 2011; Tauris et al., 2011, 2012). Specifically, their current orbital periods are orders of magnitude shorter than the Roche-lobe filling periods of their presumed AGB progenitor stars, possibly indicating an inspiral within a common envelope following unstable Roche-lobe overflow. Surprisingly, despite this different evolution, the typical measured eccentricities of these two classes are rather similar ($e\sim 10^{-6}-10^{-3}$), leading us to speculate whether they were excited by the same mechanism. Other models might face a fine-tuning challenge in trying to explain such tiny – but finite – eccentricities.
@@
To answer this question, we extended the Phinney (1992) model to AGB stars that initially fill their Roche lobe. Due to the increasing role of radiation pressure in their interiors, the $P(m_{\rm c})$ relation of these stars flattens (Rappaport et al., 1995), such that the mass of the envelope $m_{\rm env}^{\rm crit}$ – rather than that of the core $m_{\rm c}$ – becomes the dominant factor in determining the eccentricity. By running a suite of mesa simulations with $\Delta m_{\rm c}=10^{-3}\,{\rm M}_{\sun}$ intervals, we resolved the thermal pulses on the AGB and found that $m_{\rm env}^{\rm crit}\propto m_{\rm c}^{-6}$. For the most massive white dwarfs, $m_{\rm env}^{\rm crit}$ is so low that the lengthening circularization time-scale $t_{\rm circ}$ freezes the eccentricity even before the progenitor star contracts within its Roche lobe (Appendix A). Quantitatively, AGB convection yields an approximately constant $e\sim 3\times 10^{-3}$, which is almost independent of $m_{\rm c}$.
@@
We compared our theoretical prediction – which has no free parameters – to the observations, finding that almost all CO white dwarf–millisecond pulsar binaries have lower eccentricities than predicted. This indicates that orbital eccentricities are damped, rather than excited, during the subsequent common envelope inspiral phase. Quantitatively, we found an empirical relation $e\propto P^{3/2}$ between the final eccentricities and periods of white dwarfs with median masses $m_{\rm c}>0.6\,{\rm M}_{\sun}$. Presumably thanks to the almost uniform Roche-lobe overflow $e$ and $P$ initial conditions in this mass range, our final $e$$P$ relation is much tighter than Hui et al. (2018), who considered less massive CO white dwarfs as well. This observed correlation motivates future theoretical work on the eccentricity evolution during the common envelope phase (see also Dermine et al., 2013; Glanz & Perets, 2021; Szölgyén et al., 2022; Trani et al., 2022).
For the less massive CO white dwarfs – with median masses $m_{\rm c}<0.6\,{\rm M}_{\sun}$ – we also considered an alternative scenario of Case A/B stable overflow before their giant progenitors have the chance to ascend the AGB (Tauris et al., 2000; Shao & Li, 2012; Misra et al., 2020). We computed a typical binary stellar evolution track of an $m_{\rm c}\approx 0.5\,{\rm M}_{\sun}$ CO white dwarf formed through this channel, reproducing $P$ and $e$ of some of the observations without an uncertain common envelope phase. Our preliminary result motivates a more comprehensive examination of the eccentricity fluctuation and dissipation in this alternative formation path.
@@
Although we focused here on radio observations of white dwarfs around millisecond pulsars, our results may be applied to overflowing AGB stars (and their white dwarf remnants) orbiting other types of stars as well. Specifically, the Gaia mission has recently uncovered a population of white dwarf–main sequence wide binaries that could have formed similarly (Hallakoun et al., 2024; Shahaf et al., 2024; Yamaguchi et al., 2024c). Yamaguchi et al. (2024b) studied a sample of such stars with white dwarf masses $m_{\rm c}>0.6\,{\rm M}_{\sun}$, similar to our sample of massive white dwarfs in Fig. 5. Interestingly, the measured eccentricities of these stars $e\sim 10^{-2}-10^{-1}$ lie above our upper limit for stable mass transfer (Fig. 4), but they fit our empirical $e\propto P^{3/2}$ relation for the post common envelope eccentricity (Fig. 5). Another recent example is the class of self-lensing binaries found by the Kepler spacecraft, which lie relatively close to (and below) the Roche-lobe filling $P(m_{\rm c})$ line of AGB stars (Kruse & Agol, 2014; Kawahara et al., 2018; Belloni et al., 2024b; Yamaguchi et al., 2024a). The high eccentricities of barium stars present a related long-standing puzzle (Jorissen et al., 1998; Pols et al., 2003). Specifically, natal white dwarf kicks due to asymmetric mass loss at the end of the AGB phase have been proposed as a possible solution (Heyl, 2007; Izzard et al., 2010; Stone et al., 2015; El-Badry & Rix, 2018). Whether or not similar kicks affect the eccentricities of binary millisecond pulsars remains to be tested.